Simplify your results. x2=x-10
Answer by edjones(8007)

x^2=x-10
x^2-x+10=0
x^2-x+1/4=-10+1/4 complete the square
(x-1/2)^2=-39/4
x-1/2=+-sqrt(-39)/2
x=1/2+(sqrt(39)i/2), x=1/2-(sqrt(39)i/2)
.
Ed